OpenStudy (anonymous):

A truck undergoes a constant acceleration of 2.0 meters/second2 on a highway that slopes upward at an angle of 30.0° to the horizontal. What is the magnitude of the horizontal component of the acceleration? 1.0 meters/second2 1.2 meters/second2 1.7 meters/second2 2.0 meters/second2

OpenStudy (anonymous):

To find the horizontal component of the acceleration use pythagoras theorem. Where the resultant acceleration (2.0m/s2) is the hypotenuse and the horizontal component is the adjacent. Giving us the equation : \[\cos \theta = \frac{ adjacent }{ hypotenuse }\] then : \[Adjacent = hypotenuse \\cos \theta \] Where \[\theta \] is equal to the angle = 90-30= 60 So, multiply the resultant acceleration by \[cos\theta \] giving you 1m/s2

OpenStudy (anonymous):

What is the period of a satellite in a circular orbit just above the surface of the Moon? The Moon’s mass is 7.36 × 1022 kilograms and its radius is 1.738 × 106 meters. 1.94 × 104 seconds 1.30 × 104 seconds 6.50 × 103 seconds 3.24 × 103 seconds 1.62 × 103 seconds